OpenStudy (anonymous):

Write an equation of a line in point-slope form that has a slope of -2 and passes through (5, -1).

OpenStudy (anonymous):

\[\frac{y-y_1}{x-x_1}=m\]

OpenStudy (anonymous):

\[\frac{y-5}{x+1}=-2\]

OpenStudy (anonymous):

\[y-5=-2(x+1)\]

OpenStudy (anonymous):

\[y-5 = -2x-2\]

OpenStudy (anonymous):

\[y = -2x+3\]

OpenStudy (anonymous):

\[y+1 \div x -5=-2\]

OpenStudy (anonymous):

\[2x+y=9\]